OpenStudy (anonymous):

help!! 1 medal and fan The length of a rectangle is 8 1/2 inches and the width is 3 1/4 inches. What is the ratio, using whole numbers, of the length to the width? a. 17:13 b. 13:17 c. 34:13 d. 13:34

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

you need \(\frac{8\tfrac{1}{2}}{3\tfrac{1}{4}}\) rewrite the mixed numbers as improper fractions

OpenStudy (anonymous):

\[\frac{\frac{17}{2}}{\frac{13}{4}}=\frac{17}{2}\times \frac{4}{13}\]

OpenStudy (anonymous):

2.6

OpenStudy (anonymous):

you want a ratio, not a decimal

OpenStudy (anonymous):

or 34/13

OpenStudy (anonymous):

that is the one!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw